OpenStudy (anonymous):

how to find two irrational numbers between 2/3 and 4/5 need help fast!!!!!!!!!!!1

OpenStudy (anonymous):

anyone two will do?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

14/19

OpenStudy (anonymous):

9/13

OpenStudy (anonymous):

thxs man

OpenStudy (anonymous):

Btw you know the definition of an irrational number right?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

irrational nos are like root 2

OpenStudy (anonymous):

non terminating non repeating decimals

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

like 9.1231531234 bar

OpenStudy (anonymous):

no bar

OpenStudy (anonymous):

bar means it repeats but you cut it off, you rounded it

OpenStudy (anonymous):

i mean hey all I did was take the calculator punch in a couple of fractions and picked the ones between .66666666 and .8

OpenStudy (anonymous):

or a repeating decimal like 2/3 0.666666666666 bar means it keeps going forever but you cut it off

OpenStudy (anonymous):

thxs

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

irrational not rational right?

OpenStudy (anonymous):

none of the numbers you have are irrational, they are all rational

OpenStudy (anonymous):

so your answer could be something like \[\frac{2}{3}+\frac{\pi}{20}\]

OpenStudy (anonymous):

or \[\frac{2}{3}+\frac{\sqrt{2}}{20}\] or something of that type. all the numbers written above are rational numbers.

OpenStudy (anonymous):

\(\frac{2}{3}\) is rational, \(.6666\) is rational, \(\overline{.6}\) is rational, in fact \(\overline{.6}=\frac{2}{3}\)

OpenStudy (anonymous):

i guess what i am trying to say is that each number from the first answers are rational, and therefore are not right.